OpenStudy (anonymous):

A purse containing $3.20 in quarters and dimes has, in all 20 coins. Find the number of each kind of coin. help x.x

OpenStudy (anonymous):

this looks mighty familiar

OpenStudy (anonymous):

i know but i have trouble applying to algerbraic expressions. sorry if im a pain. :(

OpenStudy (anonymous):

q + d = 20 .25q + .1d = 3.2

OpenStudy (anonymous):

solve it like you did the last one exactly start with \(25x+10(20-x)=320\) as you did before right?

OpenStudy (anonymous):

Solve for d in the first equation as @satellite73 showed

OpenStudy (anonymous):

then substitute that value into the second (bottom) equation for d.

OpenStudy (anonymous):

then just as before, multiply out via the distributive law to get \[25x+200-10x=320\] combine like terms to get \[15x+200=320\] and two more steps

OpenStudy (anonymous):

@xxjennxx123 you see that this is almost identical to the last one right? only difference is the numbers, but the method and steps are the same

OpenStudy (anonymous):

yeah i see it i just have trouble applying these skills :(

OpenStudy (anonymous):

that is why you have homework and you are not being a pain, lots of people will be happy to answer

OpenStudy (anonymous):

awuh thank you :')

OpenStudy (anonymous):

pretty much when solving a linear equation you have to follow the steps remove parentheses using distributive law combine like terms add or subtract to get variable on one side, number on the other multiply or divide to get the variable by itself

OpenStudy (anonymous):

not always, but mostly

OpenStudy (anonymous):

i try to take my time with these questions, but the teacher moves on so fast. =/

OpenStudy (anonymous):

He or She have tutoring?

OpenStudy (anonymous):

You've got to take advantage of it if they do. "Too fast" is something you can discuss with them during tutoring hours.

OpenStudy (anonymous):

yeah, well it just gets worse, sorry to say they get faster and faster the higher up you go by the time you take calculus 12 they will go so fast it will break the sound barrier

OpenStudy (anonymous):

ohhh..

OpenStudy (anonymous):

Lol @satellite73 . It's not that bad @xxjennxx123. Trust me.

OpenStudy (anonymous):

You mature as you move up in math. You'll be ok. Just go to tutoring. NOW.

OpenStudy (anonymous):

trust me, it is worse than i said

OpenStudy (anonymous):

It's not. Don't do that to her! LOL!!

OpenStudy (anonymous):

ok kidding and anyway there is no such thing as calculus 12

OpenStudy (anonymous):

FACT!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

I actually thought you meant calculus in 12th grade @satellite73! Madness...calculus 12!! : )

OpenStudy (anonymous):

Anyway @xxjennxx123 seriously, use tutoring. Don't be lazy. Take care...later

OpenStudy (anonymous):

okk haha i will thanks for the encouragement. :)